OpenStudy (kitten_is_back):

Which value of y makes the equation true? y – 6.3 = –2.2

OpenStudy (lynfran):

solve for y by adding 6.3 to both sides of the equation

OpenStudy (kitten_is_back):

8.5 AND 12./6

OpenStudy (lynfran):

y-6.3=-2.2 y-6.3+6.3=-2.2+6.3 y= -2.2+6.3